OpenStudy (anonymous):

logx9 = -2 . what's the value of x?

OpenStudy (anonymous):

the log argument is 9x or x^9??

OpenStudy (anonymous):

And what's the base of the logarithm?

OpenStudy (anonymous):

the base is x

OpenStudy (anonymous):

Ohhhh. Much easier, then. So, x^(-2) = 9 which means that x = 9^(1/2) or -9^(1/2) = 1/3 or -1/3.

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

Should the negative term not be in there? I wasn't completely positive. :P

OpenStudy (anonymous):

(Pun.)

OpenStudy (anonymous):

not there... the correct way write is 9=x^-2 9=1/x^2 x^2=1/9 x=+1/3 or x=-1/3

OpenStudy (anonymous):

Yes, I just skipped a step. Sorry :P